ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 49, 90 Simelienè, Nijolè, Lithuania, Bistrickas, Vilius, Lithuania Fundamental models for several levels inventory control processes Stochastic several levels of orders inventory control along the market process is approximated by the fundamental model. Salesmen of corresponding levels in this process serve orders of their levels. Some salesmen of these levels serve orders of higher levels. They determine the prices of orders of corresponding levels. Payments for salesmen of service levels are the same and no less than their expenses. This fundamental model is employed for the approximation of analogue several levels of orders inventory control process with stochastic constraints for the storage of products.